# Bayesian spatiotemporal modelling of political violence and conflict events using discrete-time Hawkes processes

### v1

@@ 2.3 Inference
Several inferential approaches, namely Bayesian methods and maximum likelihood estimation (MLE), are used in this work. In the first instance, we adopt the Bayesian framework, due to its advantages such as the ability to directly quantify uncertainty and introduce prior knowledge from experts or previous studies directly in the inference. This model was fitted using the probabilistic programming language Stan (Stan Development Team, 2022), which implements MCMC sampling through the No U-Turn Sampler (NUTS). Prior sensitivity analysis found that the algorithm did not converge when placing uniform priors on the model parameters. Thus, weakly informative priors were placed on all model parameters such that,
  • $\mu\sim\text{Gamma}(2,2)$,
  • $\alpha\sim\text{Gamma}(2,2)$,
  • $\beta\sim\mathcal{U}(0,1)$,
  • $\sigma\sim\mathcal{IG}(5,5)$,
whereby ‘Gamma’ is the Gamma distribution, ‘$\mathcal{U}$’ the uniform distribution and ‘$\mathcal{IG}$’ the inverse Gamma distribution. While the Bayesian approach has many advantages, this comes at the expense of computational time. As such, we also consider a maximum likelihood approach where the optimisation was performed in R using the optim() function. In addition to being a less computationally intensive alternative, the MLE can also be used as a starting point for the Bayesian method.
@@ 3 Results
We first present the results obtained from fitting the Bayesian spatiotemporal model described in Sections 2.2 and 2.3 to the monthly counts of conflict events from the 2010 – 2014 ACLED data for each country and conflict type combination detailed in Section 2.1. We also propose an alternative inference scheme based on maximum likelihood estimation, and compare the modelled estimates obtained through both approaches.
In this study, the maximum excitation time $t_{\max}$ was assumed to be 3 months. The impact of this assumption in our modelling is that past events cannot affect the intensity of the process after 3 months have elapsed. As such, the model was trained on data starting from the fourth month of 2010 to mitigate edge effects and ensure that the model is not biased due to past events not being included in the intensity function calculations. To determine the timescale of temporal correlation in the data, exploratory data analysis was performed and the ARIMA model (Hyndman and Athanasopoulos, 2018) was fitted to the data in R using the auto.arima() function from the forecast package. The best model found for each country and conflict type combination indicated that all maximum lags were less than 3 months. The full analysis can be found in the supplementary materials.
3.1 Bayesian Spatiotemporal DTHP
@@
In this section, we first describe the fitting and then compare the observed event counts with the expected number of events estimated from our model. The posterior triggering kernels are also visualised and we present maps of the posterior median event counts. Lastly, out-of-sample posterior predictive checks are shown.
The Bayesian spatiotemporal DTHP with intensity function given by (3) was fitted to the data described in Section 2.1. As discussed in Section 2.3, the fitting was performed in Stan. The $\hat{\mathcal{R}}$ diagnostic (Vehtari et al., 2021) suggested that all models successfully converged. This was further validated by inspecting convergence diagnostic plots including autocorrelation functions and trace plots, and confirming that there were no divergent transitions that would indicate Stan could not find algorithm settings that allowed the posterior to be explored, while not failing. These diagnostics can be found in the supplementary materials. Furthermore, implementation code is available at https://github.com/RaihaTuiTaura/st-hawkes-thesis.
To learn about short term trends, Figure 1 shows the expected number of events for each month, estimated by the intensity function $\lambda(t,x,y)$, and compares these to the observed event counts. As a result of using the Bayesian methodology, we obtained 95% posterior intervals to represent the uncertainty around these estimated values. For each month, the event counts at each location were summed to obtain the total monthly event count. The estimated mean number of events in each month is in line with the observed data and the model is able to react quickly to recent events. Furthermore, updated estimates of conflict risk can be calculated as new data arrives by simply re-evaluating the intensity function using the estimated model parameters.
Figure 1: Observed data on month $t$ (black line) versus estimated $\lambda(t)$ (red line) and 95% posterior interval (red ribbon). Bayesian model estimated via Stan.
The posterior median excitation kernels estimated from the Bayesian model for each country are presented in Figure 2. For several conflict types Nepal has a quickly decaying temporal triggering kernel, suggesting that the level of self-excitation is highest in the first month following an event and subsides quickly thereafter. While for the remaining scenarios, the temporal kernels generally have a longer memory, meaning the influence of past events can persist as time elapses.
@@
Furthermore for most conflict types, spatial clustering of conflict events is more localised in Bangladesh and Pakistan with quickly decaying spatial triggering kernels. On the other hand, Nepal and Sri Lanka have more slowly decaying kernels for all conflict types excluding Protests and Riots, suggesting that the influence to the intensity function for events occurring in nearby locations and those further away are more similar compared to other conflict types.
@@
(a) Estimated temporal excitation kernels.
@@
(b) Estimated spatial excitation kernels. Units for horizontal axis are distance between two sets of coordinates, measured by Euclidean distance.
@@
Figure 2: Posterior median estimated excitation kernels for all countries.
Figure 3 presents a map of the median conflict risk over the 5 year observation window considered in this study. The maps show that most activity occurs in or near the capital cities for Bangladesh, Nepal and Pakistan. Pakistan also has a high concentration of protests on the southeastern border. Areas of the maps that are missing correspond to locations where no events were recorded over the 5 year period. These missing areas are particularly pronounced in Sri Lanka, which experienced a much lower volume of events over the observation window. Maps of the absolute value of the corresponding residuals for each country can be found in the supplementary materials. The residuals are generally quite low, with larger residuals corresponding to areas with a higher volume of events.
Figure 3: Posterior median number of events over 5 year observation window
@@
To assess the predictive ability of the Bayesian model, Figures 4 and 5 show the results from out-of-sample posterior predictive checks. Using 100 posterior samples, we simulated three months of data following on from the end of the 5 year observation window, which corresponds to predicting the first three months of 2015. The results were then aggregated, over both location and time, for each country and conflict type pair. Overall, the 95% interval for the predicted number of events usually encapsulates the observed number of events. However there are some larger than expected event counts that are not captured in our prediction intervals. This is likely because there is no mechanism in our model to allow for more extreme event counts, such as an over-dispersion parameter.
@@
Figure 4: Out of sample predictive checks, aggregated over 3 month prediction interval for each location and conflict type. The black dots show the observed number of events. The red bars show the 95% interval of simulated events using parameter values from 100 posterior samples.
Figure 5: Out of sample predictive checks, aggregated over all locations for each month and conflict type. The black dots show the observed number of events. The red bars show the 95% interval of simulated events using parameter values from 100 posterior samples.
3.2 Maximum likelihood approach
While the Bayesian approach presented in Section 3.1 has many advantages, as discussed in Section 2.3, maximum likelihood estimation is a less computationally expensive approach. Therefore, here we present this alternative inference method and compare the estimated model with the Bayesian approach. The point estimates obtained via maximum likelihood estimation for all countries and conflict types and the corresponding estimates from the Bayesian approach are presented in Figure 6. In most cases the maximum likelihood estimates (MLEs) lie within or are close to their respective 95% posterior credible intervals. The intensity functions estimated also align closely with the observed data, as in the Bayesian approach. The MLEs obtained and plots of the estimated intensity functions are available in the supplementary materials.
The magnitude parameter $\alpha$ indicates the level of self-excitation present in the data. For all countries, strategic developments have the lowest level of self-excitation with the lowest $\alpha$ parameters, while, protests have the highest level out of all conflict types. In Sri Lanka, protests are the only conflict type that exhibits a reasonably high level of self-excitation, since $\alpha$ is very low for all other types. Overall, Pakistan displays the highest tendency for self-exciting behaviour out of all countries, with a consistently higher or comparable $\alpha$ for all conflict types. The inverse mean of the geometric temporal triggering kernel $\beta$ is generally low, with the exception of several conflict types in Nepal. This indicates that the influence of past events on the future intensity is more sustained throughout the 3 month excitation time. The background rate $\mu$ is generally higher for protests and riots, indicating that, compared to other conflict types, more protests and riots arise independently as opposed to arising due to other events in the process. The spatial influence of these conflict types also tends to be more localised, as indicated by a small $\sigma$, the bandwidth for the spatial triggering kernel.
Figure 6: Comparison of estimated parameters for all conflict types and countries. MLE point estimates (blue) and posterior median and 95% interval (red).
Figure 7 presents a comparison of the estimated model parameters and corresponding uncertainty quantification obtained using both inference approaches considered throughout this article for a single example scenario, namely Battles in Bangladesh. To calculate the standard errors of the MLEs, we assume asymptotic normality and subsequently take the square root of the negative inverse Hessian matrix. In the Bayesian setting, standard 95% posterior credible intervals are used. It is apparent that the Bayesian approach gives much more precise estimates of parameter uncertainty. Additionally, we found in some scenarios that the Hessian matrix was singular. Thus, while the MLE approach can give reasonable point estimates, it should be used with caution when being used for uncertainty quantification. Overall, the Bayesian approach is preferable when possible, but MLEs can be a reasonable alternative when computational resources are limited provided the results are carefully scrutinised, and if the user is mainly concerned with point estimates.
Figure 7: Battles in Bangladesh. 95% confidence intervals based on MLE approach versus 95% credible intervals based on Bayesian approach.
@@ 4 Assessing conflict risk in practice
Compared to relying on historical moving averages, a key advantage of using our model is the ability to mitigate the impact of noisy data and obtain stable estimates of risk. Our approach is also more interpretable than other statistical methods based on machine learning techniques, and our Bayesian approach provides a more precise understanding of uncertainty. By incorporating the time elapsed since past events and considering the influence of spatial distance between events in a sophisticated, yet interpretable way, we can provide a more robust foundation for monitoring risk and developing early warning tools, thus informing the allocation of resources for anticipatory actions.
In this section, we demonstrate through practical examples how our statistical model provides valuable insights that enable stakeholders to make more informed decisions about how to address conflict risks. In particular, we illustrate three types of insights from our modelling approach that may be useful for decision makers: an early warning tool for monitoring current trends, visualisations of the spatial risk profile of conflict events, and short term predictions about the future risk of conflict events.
@@
The Trendfinder dashboard calculates a rolling 12-month historical average of event counts, and flags time periods where the event count is more than two standard deviations above the rolling average. To benchmark this approach against our model, we instead used the quantile function of a Poisson distribution with mean given by the estimated intensity function from our model to determine time periods with unusually high activity. In particular, the estimated intensity function was calculated using 100 posterior samples of the model parameters. Then, for each posterior estimate of the intensity function, we obtained the median number of events and corresponding 95% interval of event counts at each month using the quantile function. Rolling 12-month averages were then taken of the median of these quantiles at each month. Time periods were then flagged if the observed number of events exceeded the rolling average of the 97.5% quantile.
@@
Figures 8 and 9 present examples of early warning tools respectively based on our modelled risk estimates, and a naive approach using simple historical moving averages over the same time period, inspired by the ACLED Trendfinder dashboard. The naive approach has very over-confident and volatile uncertainty bounds, whereas our approach provides wider and more stable bounds. As a result, the naive approach is highly sensitive and flags many more months as having unusually high activity. Many of the event counts in months flagged using the naive approach are within the reasonable range of values expected from our model based on the estimated intensity function. Furthermore, in Sri Lanka a single event from either the battles or explosions/remote violence categories is flagged using the naive approach. Our approach is also more robust to outliers, since these events do not directly contribute to the calculation of the rolling average. The 97.5% trigger threshold was chosen arbitrarily in this example, but this threshold percentage can also be adapted to suit the risk appetite of the user.
@@
Figure 8: Hawkes model. Grey bars: no flag. Blue bars: flagged months (observed counts greater than 97.5% rolling 12-month average. Red line and ribbon: 12-month rolling average of the median 50%, 2.5% and 97.5% quantiles from the Poisson distribution, calculated from 100 posterior samples of the estimated intensity function.
Figure 9: Historical averages (based on ACLED Trendfinder dashboard). Grey bars: no flag. Blue bars: flagged months (observed counts more than two standard deviations above the rolling 12-month historical average. Red line and ribbon: 12-month rolling historical average plus or minus two standard deviations.
@@
The visualisation of the spatial risk profile for conflict events is also an important aspect of conflict risk monitoring. As an example, Figure 10 presents the median number of events expected during December 2014, the last month of the observation window considered in these analyses. Our modelling operates at a finer spatial scale than current methods employed by organisations such as ACLED. This granularity takes into account the unique attributes of each administrative region. Furthermore, the estimates can easily be aggregated up to obtain more accurate estimates of a coarser spatial resolution, with corresponding estimates of uncertainty.
Figure 10: Risk of conflict events at December 2014 (i.e. the end of the observation window)
@@
Another notable advantage of adopting a statistical model for analysing conflict risk is the ability to generate short-term predictions of future events. In addition to estimating the level of future risk, our Bayesian approach also produces estimates of uncertainty surrounding the predicted number of events. In Section 3.1, we used 100 posterior samples to predict the expected number of events in the first 3 months of 2015, namely 3 months from the end of our modelling observation window. Figures 11 to 13 present the 2.5%, 50% and 97.5% percentiles of the cumulative number of events predicted over these 3 months. Across most regions, the projected risk level for these three months is relatively low. However, there are numerous hotspots, particularly in Pakistan, where even the predictions for the lowest percentile suggest a substantial number of events. These spatial predictions, along with their uncertainty intervals, provide valuable insights into the distribution and intensity of anticipated conflict events, enabling policymakers and stakeholders to prioritise resources and interventions accordingly.
Figure 11: Out-of-sample 3-month-ahead predictions. 2.5% percentile of predicted events.
Figure 12: Out-of-sample 3-month-ahead predictions. 50% percentile of predicted events.
Figure 13: Out-of-sample 3-month-ahead predictions. 97.5% percentile of predicted events.

### v2

@@ 2.3 Spatiotemporal discrete-time Hawkes process
In this work we adopt a Bayesian framework due to its advantages, such as the ability to directly quantify uncertainty and introduce prior knowledge from experts or previous studies directly in the inference. This model was fitted using the probabilistic programming language Stan (Stan Development Team, 2022), which implements MCMC sampling through the No U-Turn Sampler (NUTS).
We modelled the spatially varying baseline rate parameters $\mu_{x,y}$ using the Intrinsic Conditional Auto-Regressive (ICAR) model (Besag, 1974). To simplify notation, we will simply write $\mu_{r}$ to refer to the region-specific baseline rate parameter for region $r$ with centroid $(x,y)\in\mathcal{S}$. We can then define the ICAR model based on Morris et al. (2019) such that each $\mu_{r}$ is distributed according to a normal distribution with mean equal to the average of its neighbours and such that,
$p(\mu_{r}|\mu_{j\in N(l)})=\mathcal{N}\bigg(\frac{\sum_{j\in N(l)}\mu_{j}}{d_{r}},\frac{s_{r}^{2}}{d_{r}}\bigg)$ (9)
where $d_{r}$ is the number of neighbours of region $r$, $N(r)$ is the set of its neighbours, and $s_{r}^{2}$ is an unknown variance. The joint distribution of $M=(\mu_{r},\ r\in\{1,\dots,|\mathcal{S}|\})$ is then,
$M\sim\mathcal{N}(0,Q^{-1})$ (10)
where $Q$ is the precision matrix. Morris et al. (2019) defines $Q$ in terms of a diagonal matrix $D$ containing the number of neighbours in each region, and an adjacency matrix $A$ that identifies neighbouring regions. The precision matrix can then be written,
Q=DA.Q=D-A. (11)
For the non-spatially varying parameters, we considered two prior specifications: weakly informative priors and priors centered around the maximum likelihood estimates (MLEs). We found that the algorithm often did not converge when placing weakly informative priors on the non-spatial model parameters due to difficulties in estimating the correlation structure of the spatially varying baseline parameters. We therefore specify the following priors centered around the MLEs,
$\mu_{0}$ $\sim\mathcal{N}(\hat{\mu_{0}},0.1)$
$\alpha$ $\sim\log\mathcal{N}(m_{\hat{\alpha}},0.1)$
$\beta$ $\sim\text{Beta}(a_{\hat{\beta}},b_{\hat{\beta}}),\ 0<\beta<1$
$\sigma$ $\sim\log\mathcal{N}(m_{\hat{\sigma}},0.1)$
$\pi$ $\sim\text{Beta}(a_{\hat{\pi_{0}}},b_{\hat{\pi_{0}}}),\ 0<\pi<1$ (12)
where ˆ indicates the MLEs which were calculated in R using the optim() function and $m$, $a$ and $b$ are constants calculated to ensure the mean of the prior is equal to the MLE. Further details of this prior sensitivity analysis are provided in the supplementary material.
@@ 3 Results
In this section we present the results obtained from fitting the Bayesian spatiotemporal models described in Section 2.3 to the weekly counts of conflict events from the 2010 – 2014 ACLED data for each country and conflict type combination detailed in Section 2.1. We first describe the fitting process, in which we compare two specifications for the spatiotemporal DTHP and, for each scenario, select the best model to use in downstream analysis. We then compare the observed event counts with the expected number of events estimated from our model. The posterior distributions for the model parameters and triggering kernels are also visualised. Lastly, out-of-sample posterior predictive checks are shown. The code to implement these models is available at https://github.com/RaihaTuiTaura/st-hawkes-thesis.
As discussed in Section 2.3, we considered two specifications for the form of the spatiotemporal DTHP. We refer to the model defined in (3) with only temporal self-excitation as “nospatialSE” and the model in 4 with self-excitation both temporally and spatially as “ST_SE”. To choose an appropriate model form, we consider the following criteria. As a first pass, the algorithm must not show signs of non-convergence. This was assessed based on diagnostic criteria such as the $\hat{\mathcal{R}}$ (Vehtari et al., 2021), diagnostic plots including autocorrelation functions and trace plots, and by confirming that there were no divergent transitions that would indicate that Stan could not find algorithm settings that allowed the posterior to be explored without failure. Only two scenarios failed this criteria, namely Battles and Protests in Nepal under the “nospatialSE” model, so for these two scenarios we use the “ST_SE” model. These diagnostics can be found in the supplementary materials.
If satisfied that the algorithm has likely converged, we consider the following goodness-of-fit criteria for model comparison to consider various aspects of the data:
  1. 1.
    Out-of-sample coverage: proportion of observations where the observed event counts fall within the 95% posterior interval based on 100 out-of-sample predictions.
  2. 2.
    In-sample coverage: proportion of observations where the observed event counts fall within the median 95% quantile interval (2.5% - 97.5% quantiles) calculated based on 100 posterior samples.
  3. 3.
    In-sample residuals (non-zero; median): sum of the absolute value of the difference between the observed event count and the median of the 50% quantile calculated based on 100 posterior samples.
  4. 4.
    In-sample residuals (non-zero; upper quantile): sum of the absolute value of the difference between the observed event count and the median of the 97.5% quantile calculated based on 100 posterior samples.
  5. 5.
    In-sample residuals (zeros): sum of the absolute value of the difference between the observed number of zero event counts and the proportion in which the median of the 50% quantile is equal to 0 calculated based on 100 posterior samples.
  6. 6.
    BIC: median BIC with likelihood calculated based on 100 posterior samples.
If the temporally and spatially self-exciting model (“ST_SE”) had better performance in greater than 50% of these goodness-of-fit criteria, we select this model; otherwise, we resort to the more parsimonious model that only includes a temporal self-excitation mechanism (“nospatialSE”). Table 1 presents the chosen model for each scenario.
Conflict type Bangladesh Nepal Pakistan Sri Lanka
Battles ST_SE ST_SE nospatialSE ST_SE
Explosions/Remote violence nospatialSE ST_SE nospatialSE ST_SE
Protests ST_SE ST_SE nospatialSE nospatialSE
Riots nospatialSE ST_SE nospatialSE nospatialSE
Strategic developments nospatialSE ST_SE nospatialSE nospatialSE
Violence against civilians nospatialSE ST_SE nospatialSE nospatialSE
Table 1: Choice of spatiotemporal DTHP for each scenario
The maximum excitation time $t_{\max}$ was assumed to be 12 months. The impact of this assumption in our modelling is that past events cannot affect the intensity of the process after 12 months have elapsed. As such, the model was trained on data starting from the beginning of 2011 to mitigate edge effects and ensure that the model is not biased due to past events not being included in the intensity function calculations.
To learn about short term trends, Figure 5 shows the modelled event counts for a single scenario, namely Protests in Pakistan and compares these to the observed event counts. Our Bayesian approach also allows us to characterise uncertainty around our estimates. We obtained the median 50% quantile and 95% uncertainty intervals for the estimated event counts by calculating, for 100 posterior samples, the 2.5%, 50% and 97.5% quantiles of event counts arising from the estimated Poisson distributions.
We see in our example for Protests in Pakistan that the median 50% quantile and the corresponding uncertainty intervals around the counts are closely aligned to the observed counts within the individual regions and the model is able to react quickly to recent events. This is generally consistent for all countries and conflict type scenarios that we considered, as can be seen in the supplementary material. We note that, for scenarios with high zero-inflation, the median of the estimated 50% quantiles is often also 0, however the uncertainty intervals generally encompass the observed event counts.
Figure 5: Protests in Pakistan (unaggregated). Observed data on week $t$ (black bars) versus median of the estimated 50% quantile (red line) and median 95% interval of event counts calculated based on 100 posterior samples (red ribbon).
We now consider aggregating over administrative regions such that, for each week and region, we calculate and sum the 5%, 50% and 95% quantiles from 100 posterior samples over all regions to obtain weekly totals, from which we then calculate the median for each quantile. Figure 6 presents the median 90% quantile intervals around these aggregated event counts and the corresponding median of the 50% quantile, compared to the observed aggregated event counts for all scenarios. The estimated median number of events in each week is in line with the observed data. We considered tighter median quantile intervals of 90% rather than 95% since aggregating across administrative regions is expected to reduce uncertainty. Although these median quantile intervals of event counts appear to overestimate the upper bound in some scenarios, this is largely an artefact of the aggregation process. In cases where many regions experience no events, the upper bound is often a small, non-zero integer, and the accumulation of such low-count regions naturally inflates the resulting uncertainty bounds.
Figure 6: All scenarios (aggregated). Observed data on week $t$ (black bars) versus median of the estimated 50% quantile (red line) and median 95% interval of event counts calculated based on 100 posterior samples (red ribbon).
Figure 7 shows $\exp(\mu_{0}+\mu_{l})$, namely the median estimated spatially varying baseline rate, for each region $l$ across conflict types. This rate reflects the level of conflict activity expected to arise independently of other events. Regions in Sri Lanka generally have a low level of baseline risk of conflict events across all conflict types. Bangladesh has higher baseline risk of Protests, Riots and Violence against civilians compared to other event types. Nepal has higher risk of Protests and Riots, and Pakistan has higher risk of Battles in the northwest region and Protests and Riots in the southeastern regions. In practice these estimates can provide actors in the humanitarian sector with a useful tool to discern the level of activity to be expected independently of any other events and the corresponding uncertainty in these estimates.
Figure 7: Median estimated baseline rates. Note that while displayed on a single map models for each country were estimated individually. Black crosses indicate capital cities (including Karachi in Pakistan as the biggest city and economic capital).
Figure 8 shows the estimated distribution of non-spatially varying model parameters. As mentioned in Section 2.3, the zero-inflation factor in our model means that the magnitude parameter $\alpha$ does not directly correspond to the expected number of offspring overall; rather, it indicates the level of self-excitation within the non-structural zero portion of the likelihood and should be interpreted in conjunction with the probability of structural zeros $\pi$. There are a number of scenarios for which the estimated $\alpha$ is essentially zero, particularly in Sri Lanka, indicating that a baseline only model may be sufficient for these scenarios. The scenarios for which this was the case fall into two categories: very few events occurring, or events occurring reasonably frequently but for which the baseline rate alone fits the data well. Large values of $\alpha$ typically correspond to high $\pi$, as seen often with Explosions/remote violence and Strategic developments, indicating that there are many zero counts but events that do occur are highly excitatory. There is a range in values of the inverse mean of the geometric temporal triggering kernel $\beta$, ranging from very low, which indicates that the influence of past events on the future intensity is more sustained throughout the 12 month excitation time, to higher values indicating that the influence drops off more sharply as time passes. The global baseline rate $\mu_{0}$ is generally higher for protests and riots indicating that, compared to other conflict types, more protests and riots arise independently rather than as a result of self-excitation mechanisms. For scenarios in which the “ST_SE” model was selected, the bandwidth for the spatial triggering kernel $\sigma$ was also estimated. In several cases $\sigma$ is very small, indicating that these conflict types also tend to be more localised; however, particularly in Nepal, we see larger $\sigma$ values which indicates a broader reach.
Figure 8: Distribution of estimated parameters
The posterior median excitation kernels and corresponding 95% intervals estimated from 100 posterior samples for each country and conflict type are presented in Figure 9. In Figure 9(a), we multiplied the temporal excitation kernels by $\alpha$ to reflect the effective level of excitation from a single past event. The temporal excitation kernels for Nepal and Sri Lanka generally decay very quickly for all conflict types. On the contrary, conflicts in Bangladesh and Pakistan generally have longer excitation memories, meaning the influence of past events can persist for longer as time elapses. This is particularly true for the conflict types Battles, Explosions/remote violence and Strategic developments.
@@
For scenarios in which the spatial triggering kernel was estimated, we notice that spatial clustering of conflict events is more localised in Bangladesh, characterised by quickly decaying spatial triggering kernels. We also observe this pattern for Battles across all countries. On the other hand, for most conflict types in Nepal and for Explosions/remote violence in Sri Lanka we see more slowly decaying kernels, suggesting that the influence on the intensity function for events occurring in nearby locations and those further away is more similar compared to other conflict types.
@@
(a) Estimated posterior distribution of temporal excitation kernels multiplied by $\alpha$.
@@
(b) Estimated posterior distribution of spatial excitation kernels for scenarios using the “ST_SE” model. Units for horizontal axis are distance between two sets of coordinates, measured by Euclidean distance.
@@
Figure 9: Posterior distribution of estimated excitation kernels for all countries and corresponding 95% posterior interval.
@@
To assess the predictive ability of our model, Figure 10 shows the results from out-of-sample posterior predictive checks. Using 100 posterior samples, we simulated three months of data following from the end of the five year observation window, which corresponds to predicting the first three months of 2015. The results were then summed over locations to obtain event counts for each week for each country and conflict type pair. Overall, the 95% interval for the predicted number of events usually encapsulates the observed number of events. However there are some larger than expected event counts that are not captured in our prediction intervals, particularly in Bangladesh for Riots, Protests and Explosions/Remote Violence. This is due to the testing period coinciding with a substantial surge in events associated with the 2015 Bangladeshi political crisis (International Crisis Group, 2015). The 12 months preceding the test period exhibited low levels of conflict activity, meaning that the contribution from the self-exciting component of the intensity function was based on a period of low event counts. Consequently, the model was unable to reproduce the excess events observed during the crisis, as out-of-sample data were not used in generating the forecasts.
@@
Figure 10: Out of sample predictive checks (aggregated over 3 month prediction interval for each location and conflict type). The black dots show the observed number of events. The red bars show the 95% interval of simulated events using parameter values from 100 posterior samples.
@@ 4 Assessing conflict risk in practice
Compared to relying on historical moving averages, a key advantage of using our model is the ability to mitigate the impact of noisy data and obtain stable estimates of risk. Our approach is also more interpretable than other statistical methods based on machine learning techniques, and our Bayesian approach provides a more precise understanding of uncertainty. By incorporating the time elapsed since past events and when appropriate considering the influence of distance between events in a sophisticated, yet interpretable way, we can provide a more robust foundation for monitoring risk and developing early warning tools, thus informing the allocation of resources for anticipatory actions.
In this section, we demonstrate through practical examples how our statistical model provides valuable insights that enable stakeholders to make more informed decisions about how to address conflict risks. In particular, we illustrate three types of insights from our modelling approach that may be useful for decision makers: an early warning tool for monitoring trends, visualisations of the spatial risk profile of conflict events, and short term predictions about the future risk of conflict events.
@@
The Trendfinder dashboard calculates a rolling 12-month historical average of event counts, and flags time periods where the event count is more than two standard deviations above the rolling average. To benchmark this approach against our model, we instead used the quantile function of a Poisson distribution with mean given by the estimated intensity function from our model to determine time periods with unusually high activity. In particular, the estimated intensity function was calculated using 100 posterior samples of the model parameters. Then, for each posterior sample and corresponding estimate of the intensity function, we obtained the 95% interval of event counts at each week using the quantile function. Rolling 12-month averages were then taken of the median of these quantiles at each week. Time periods were then flagged if the observed number of events exceeded the rolling average of the 97.5% quantile.
@@
Figure 11 presents an example of this early warning tool, comparing our modelled risk estimates and a naive approach using simple historical moving averages over the same time period, inspired by the ACLED Trendfinder dashboard. We present here the result for Protests in Pakistan as an example, and the figures for the remaining scenarios are available in the supplementary materials. In general, we find that the upper bounds used as the threshold in the naive approach are highly volatile and sensitive to outliers, often resulting in large, abrupt increases and reductions in the threshold. Our approach has much more consistent intervals and is also more robust to outliers, since these events do not directly contribute to the calculation of the rolling average. For regions with lower event counts, the naive model often identifies weeks with just a single event as having unusually high activity, whereas our model often considers this to fall within the reasonable range of values expected based on the estimated intensity function. For regions with higher event counts, under the naive model, outliers can strongly influence the 12-month rolling average and inflate the threshold for a trigger, whereas our model may still identify such events as anomalous. The 97.5% trigger threshold was chosen arbitrarily in this example, but this threshold percentage can also be adapted to suit the risk appetite of the user.
Figure 11: Comparison of early warning triggers for Protests in Pakistan.
Grey bars: no flag.
Blue: flagged via naive model. Observed counts greater than two standard deviations above 12-month rolling average.
@@
Red: flagged via proposed Hawkes model. Observed counts greater than 12-month rolling average of the median 97.5% quantiles from the Poisson distribution, calculated from 100 posterior samples.
Purple: flagged in both approaches.
@@
The visualisation of the spatial risk profile for conflict events is also an important aspect of conflict risk monitoring. As an example, Figure 12 presents the median of the 50% quantile, calculated from 100 posterior samples, in the last week of December in 2014, namely the last week of the observation window considered in these analyses. Our modelling can operate at a finer spatial scale than current methods produced by organisations such as ACLED. This granularity can then take into account the unique attributes of each administrative region. Furthermore, the estimates can easily be aggregated up as required to obtain more accurate estimates of a coarser spatial resolution, with corresponding estimates of uncertainty.
Figure 12: Median of the 50% quantile, based on 100 posterior samples, at December 2014 (i.e. the end of the observation window). Black crosses indicate capital cities (including Karachi in Pakistan as the biggest city and economic capital).
@@
Another notable advantage of adopting a statistical model for analysing conflict risk is the ability to generate short-term predictions of future events. In addition to estimating the level of future risk, our Bayesian approach also produces estimates of uncertainty surrounding the predicted number of events. In Section 3, we used 100 posterior samples to predict the expected number of events in the first 3 months of 2015, namely 3 months from the end of our modelling observation window. Figure 13 presents the 2.5%, 50% and 97.5% quantiles of the cumulative number of events predicted over these 3 months. Across most regions, the projected risk level for these three months is relatively low. However, there are numerous hotspots, particularly in Pakistan and Bangladesh. These spatial predictions, along with their uncertainty intervals, provide valuable insights into the distribution and intensity of anticipated conflict events, enabling policymakers and stakeholders to prioritise resources and interventions accordingly.
Figure 13: Out-of-sample 3-month-ahead predictions.
First row: observed cumulative event counts.
Second row: median of cumulative event counts.
Third row: 2.5% quantile of cumulative event counts.
Fourth row: 97.5% quantile of cumulative event counts.
Black crosses indicate capital cities (including Karachi in Pakistan as the biggest city and economic capital).